22 At what rate compounded semi – annually will RM 2000 become RM3500 in five years?
Answer #1

Here we will use the following formula:

FV = PV * (1 + r%)n

where, FV = Future value = 3500, PV = Present value = 2000, r = rate of interest , n= time period = 5 *2 = 10 semi annual years.

now, putting theses values in the above equation, we get,

3500 = 2000 * (1 + r)10

3500 / 2000 = (1 + r)10

1.75 = (1 + r)10

(1.75)1/10 = 1 + r

(1.75)0.1 = 1 + r

1.05755705034 = 1 + r

r = 1.05755705034 - 1

r = 0.05755705034 or 5.76%

This is semi annual rate.

Annual rate = 5.76 * 2 = 11.51%

So, required rate is 11.51% compounded semi annually.